OpenStudy (anonymous):

What is the value of x, when x2 + 5 = 21?

OpenStudy (anonymous):

x^2+5=21 -> x^2=16. So, x=4 or x=-4

OpenStudy (unklerhaukus):

take away five from both sides then take the square root

OpenStudy (anonymous):

well the first thing to do is isolate x x^2 + 5 = 21 x^2 = 21 - 5 x^2 = 16 Now to take out the ^2 you find the square root x^2 = 16 x = sqrt(16 x = 4 or x = -4; since it will result the same

OpenStudy (anonymous):

Thankk youu : )